# Discrete-to-continuum limits of semilinear stochastic evolution equations in Banach spaces

### v1

@@ 3.4. Proof of convergence
In order to prove parts ac, we will respectively apply Theorems 5.46.6 and Corollary 6.10, which are the rigorous counterparts of the discrete-to-continuum result Theorem 1.1 in the respective settings.
@@
b Now we need Assumption 3.6(i) in order for $([\mathcal{L}_{n}^{\tau,\kappa}]^{s})_{n\in\overline{\mathbb{N}}}$ to be uniformly sectorial of angle less than $\frac{1}{2}\pi$ on $(L^{q}(\mathcal{M}_{n}))_{n\in\overline{\mathbb{N}}}$ for all $q\in[2,\infty)$. Letting $\delta>0$ and $\beta\in(\frac{m+4+\delta}{2s},\frac{1}{2})$ be as in Assumption 3.5(ii), it follows from Theorem 3.13b that $\widetilde{R}_{n}^{\beta^{\prime}}\to R_{\infty}^{\beta^{\prime}}$ in $L^{0}(\widetilde{\Omega};\mathscr{L}(L^{2}(\mathcal{M});L^{\infty}(\mathcal{M})))$ as $n\to\infty$, for all $\beta^{\prime}\geq\beta$, under Assumptions 3.4(ii) and 3.8. In particular, we have $\widetilde{R}_{n}^{\beta}\to R_{\infty}^{\beta}$ in $L^{0}(\widetilde{\Omega};\gamma(L^{2}(\mathcal{M});L^{q}(\mathcal{M})))$ for all $q\in[1,\infty)$ by [49, Corollary 9.3.3], and $\widetilde{R}_{n}\to R_{\infty}$ in $L^{0}(\widetilde{\Omega};\mathscr{L}(L^{\infty}(\mathcal{M})))$. This shows (A2-B) and (A3-B). By Lemma 3.14a, we have (A4-B) with $\theta=\frac{4}{q}\beta$. Thus, choosing $q>\frac{4\beta}{1-2\beta}$ yields $\theta+2\beta<1$. Conditions (F1-B) and (F2-B) follow similarly to part a.
c As in part b, we need to verify conditions (A1-B)(A4-B), now with contractive semigroups $(S_{n}(t))_{t\geq 0}$, i.e., Assumption 3.6(ii). For $s=1$, $(S_{\infty}(t))_{t\geq 0}$ is $L^{\infty}(\mathbb{T})$-contractive since the $L^{1}(\mathbb{T})$-norm of its heat kernel coincides with the $L^{1}(\mathbb{R})$-norm of the Gauss–Weierstrass kernel, which is equal to $1$. For $n<\infty$, the $L^{\infty}(\mathcal{M}_{n})$-contractivity of $S_{n}(t)=e^{-tA_{n}}$ is equivalent to $A_{n}$ being diagonally dominant with positive diagonal by [64, Lemma 6.1], which holds for Laplacian matrices. Since these assertions can be extended to all $s\in(0,1]$ by subordination, see for instance [50, Theorem 15.2.17], we indeed find that Assumption 3.6(ii) holds. Thus, we can proceed to argue as in b, using Theorem 3.13c and Lemma 3.14b for an arbitrary $\beta\in(\frac{1}{4s},\frac{1}{2})$, to obtain (A1-B)(A4-B) with $\widetilde{M}_{S}=1$ and $\theta+2\beta<1$ for $q\in[2,\infty)$ large enough.
@@ 4. Infinite-dimensional Ornstein–Uhlenbeck process
Here, $A_{n}\colon\mathsf{D}(A_{n})\subseteq E_{n}\to E_{n}$ is a linear operator and $T\in(0,\infty)$ is a time horizon. Moreover, we take $W_{n}\coloneqq\Pi_{n}W_{\infty}$, where $(W_{\infty}(t))_{t\geq 0}$ denotes a cylindrical Wiener process on $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$ taking values in a separable Hilbert space $(H_{\infty},\langle\,\cdot\,,\,\cdot\,\rangle_{H_{\infty}})$ and the operator $\Pi_{n}\in\mathscr{L}(H_{\infty};H_{n})$ is as in assumption (A1) below. Thus, the formal time derivative $\dot{W}_{\infty}$ of $W_{\infty}$ represents space–time Gaussian white noise and $(W_{n})_{t\geq 0}$ is an $H_{n}$-valued $Q_{n}$-cylindrical Wiener process colored in space by $Q_{n}\coloneqq\Pi_{n}\Pi_{n}^{*}\in\mathscr{L}(H_{n})$.
@@
    The operators $(A_{n})_{n\in\overline{\mathbb{N}}}$ on $(E_{n})_{n\in\overline{\mathbb{N}}}$ are uniformly sectorial of angle less than $\frac{1}{2}\pi$ (see Appendix C for the definition of this concept). In particular, their negatives generate bounded analytic $C_{0}$-semigroups $(S_{n}(t))_{t\geq 0}\subseteq\mathscr{L}(E_{n})$ which satisfy Assumption 2.2 with $w=0$. Moreover, there exists a $\beta\in[0,\frac{1}{2})$ such that $R_{n}^{\beta}\coloneqq(\operatorname{Id}_{n}+A_{n})^{-\beta}\in\gamma(H_{n};E_{n})$ for every $E_{n}$.
@@
As remarked in Section 2.2, we mainly have applications in mind where the problem corresponding to $n=\infty$ is interpreted as a spatiotemporal stochastic partial differential equation, whose solution is also known as an infinite-dimensional Ornstein–Uhlenbeck process, and solutions to (4.1) for $n\in\mathbb{N}$ are spatially discretized approximations. Therefore, it is natural to ask whether we can identify the right mode of convergence of the operators $(A_{n})_{n\in\mathbb{N}}$ to $A_{\infty}$ as $n\to\infty$ to ensure the convergence of the processes $(W_{A_{n}})_{n\in\mathbb{N}}$ to $W_{A_{\infty}}$.
@@
Before stating any discrete-to-continuum results, let us introduce some convenient notation for this goal. Using the operators $(\Lambda_{n})_{n\in\mathbb{N}}$ and $(\Pi_{n})_{n\in\mathbb{N}}$, we can take a mapping which has $E_{n}$ as its domain or state space, and turn it into an analogous mapping from or to $E_{\infty}$. For instance, we define the $E_{\infty}$-valued processes $\widetilde{W}_{A_{n}}\coloneqq\Lambda_{n}W_{A_{n}}$, as well as the operators $\widetilde{A}_{n}^{-\alpha}\coloneqq\Lambda_{n}A_{n}^{-\alpha}\Pi_{n}$ and $\widetilde{S}_{n}(t)\coloneqq\Lambda_{n}S_{n}(t)\Pi_{n}$ in $\mathscr{L}(E_{\infty})$ for $\alpha,t\in[0,\infty)$. Now we can formulate our notion of the convergence $A_{n}\to A_{\infty}$ as $n\to\infty$ as follows:
@@
    For every $x\in E_{\infty}$, we have $\widetilde{R}_{n}x\to R_{\infty}x$ in $E_{\infty}$ as $n\to\infty$. Moreover, given $\beta\in[0,\frac{1}{2})$ as in (A2), we have $\widetilde{R}_{n}^{\beta}\to R_{\infty}^{\beta}$ in $\gamma(H_{\infty};E_{\infty})$.
@@
Arguing as in the proof of Proposition B.3 and using ii, it follows that we only need to prove the claim for all $R$ belonging to some dense subset $D$ of $L^{2}(a,b;\gamma(H;E))$. Note that every $R\in L^{2}(a,b;\gamma(H;E))$ can be approximated by a step function $\sum_{j=1}^{N}\mathbf{1}_{(a_{j},b_{j}^{\prime})}\otimes T_{j}$ with $a<a_{j}^{\prime}<b_{j}^{\prime}<b$ and $T_{j}\in\gamma(H;E)$, and by definition of $\gamma(H;E)$ the latter can be approximated by finite-rank operators. By linearity, it thus suffices to prove the statement for $R$ of the form
@@
Substituting this representation, using (2.5) and i, we find as $n\to\infty$:
@@
condition ii is verified by applying (C.3) to $\|M_{n}(t)\|_{\mathscr{L}(E_{\infty})}$ combined with Assumptions (A1) and (A2) as in (4.16), and hypothesis i holds by Lemma C.1. ∎
@@
The proof is based on the Da Prato–Kwapień–Zabczyk factorization method, first formulated in [19] and for Hilbert spaces (see also [20, Section 5.3]), and later extended to UMD-type-2 Banach spaces in [9, Theorem 3.2]. The general idea is to express the process $W_{A_{n}}$ as the ‘product’ $\mathfrak{I}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}W^{\frac{1}{2}+\beta^{\prime}}_{A_{n}}$ of a fractional parabolic integral operator $\mathfrak{I}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}$ as in Appendix B and auxiliary process $W^{\frac{1}{2}+\beta^{\prime}}_{A_{n}}$ as in (4.8), and using the smoothing properties of the former.
@@ 5. Approximation of semilinear stochastic evolution equations with additive cylindrical Wiener noise
In Section 5.1 we will assume in particular that the drifts $(F_{n})_{n\in\overline{\mathbb{N}}}$ are uniformly globally Lipschitz and of linear growth to obtain unique global solutions $(X_{n}(t))_{t\in[0,T]}$, whose lifted counterparts $\widetilde{X}_{n}$ converge to $X_{\infty}$ in $L^{p}(\Omega;C([0,T];E_{\infty}))$ as $n\to\infty$, where $p\in[1,\infty)$ is the stochastic integrability of the initial data. These assumptions are relaxed in Section 5.2, where we suppose that the drifts are uniformly locally Lipschitz and uniformly bounded near zero. In general, this comes at the cost of obtaining merely local solutions, converging in a weaker norm. However, if one can show independently that the solutions are global and the $L^{p}(\Omega;C([0,T];E_{n}))$-norms of $X_{n}$ are uniformly bounded in $n\in\mathbb{N}$, then we recover the stronger sense of convergence.
@@ 5.1. Globally Lipschitz drifts of linear growth
$\widetilde{\Phi}_{n,T}(u)=\widetilde{S}_{n}\otimes\xi_{n}+\widetilde{S}_{n}*\widetilde{F}_{n}(\,\cdot\,,u)+\widetilde{W}_{A_{n}}$ (5.15)
@@
$\|X_{\infty}-\widetilde{X}_{n}\|_{L^{p}(\Omega;C([0,\frac{3}{2}T_{0}];E_{\infty}))}$ $\leq\|X_{\infty}-\widetilde{X}_{n}\|_{L^{p}(\Omega;C([0,T_{0}];E_{\infty}))}$ (5.34)
@@ 6.1. Setting and convergence for globally Lipschitz drifts
Fix a $\beta^{\prime}\in(\beta,\frac{1}{2})$ such that $\theta+2\beta^{\prime}<1$, where $\beta\in[0,\frac{1}{2})$ is as in (A2). Without loss of generality, way may assume that $p\in(2,\infty)$ is so large that, in fact, $\theta+2\beta^{\prime}<1-\frac{2}{p}$. As in the proof of Proposition 4.5, we see that $W_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}(\omega,\,\cdot\,)$, where $W_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}$ is the auxiliary process defined by (4.8), belongs to $L^{p}(0,T;E_{n})$ for $\mathbb{P}$-a.e. $\omega\in\Omega$, hence $\mathfrak{I}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}W_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}(\omega,\,\cdot\,)\in C([0,T];B_{n})$ by applying Proposition B.1b with $E\coloneqq E_{n}$, $F\coloneqq B_{n}$ and $\alpha\coloneqq\theta$ from assumption (A4-B). Moreover, one finds that the process $\mathfrak{I}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}W_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}$ is a continuous modification of $W_{A_{n}}$, belonging to $L^{p}(\Omega;C([0,T];B_{n}))$. It now suffices to establish the following, as $n\to\infty$:

### v2

@@ 3.4. Proof of convergence
To prove parts ac, we will apply Theorems 5.46.6 and Corollary 6.10, respectively, which are the rigorous counterparts of the discrete-to-continuum Theorem 1.1 in the respective settings.
@@
b Now we need Assumption 3.6(i) in order for $([\mathcal{L}_{n}^{\tau,\kappa}]^{s})_{n\in\overline{\mathbb{N}}}$ to be uniformly sectorial of angle less than $\frac{1}{2}\pi$ on $(L^{q}(\mathcal{M}_{n}))_{n\in\overline{\mathbb{N}}}$ for all $q\in[2,\infty)$. Letting $\delta>0$ and $\beta\in(\frac{m+4+\delta}{2s},\frac{1}{2})$ be as in Assumption 3.5(ii), it follows from Theorem 3.14b that $\widetilde{R}_{n}^{\beta^{\prime}}\to R_{\infty}^{\beta^{\prime}}$ in $L^{0}(\widetilde{\Omega};\mathscr{L}(L^{2}(\mathcal{M});L^{\infty}(\mathcal{M})))$ as $n\to\infty$, for all $\beta^{\prime}\geq\beta$, under Assumptions 3.4(ii) and 3.9. In particular, we have $\widetilde{R}_{n}^{\beta}\to R_{\infty}^{\beta}$ in $L^{0}(\widetilde{\Omega};\gamma(L^{2}(\mathcal{M});L^{q}(\mathcal{M})))$ for all $q\in[1,\infty)$ by [49, Corollary 9.3.3], and $\widetilde{R}_{n}\to R_{\infty}$ in $L^{0}(\widetilde{\Omega};\mathscr{L}(L^{\infty}(\mathcal{M})))$. This shows (A2-B) and (A3-B). By Lemma 3.15a, we have (A4-B) with $\theta=\frac{4}{q}\beta$. Thus, choosing $q>\frac{4\beta}{1-2\beta}$ yields $\theta+2\beta<1$. Conditions (F1-B) and (F2-B) follow similarly to part a.
c As in part b, we need to verify conditions (A1-B)(A4-B), now with contractive semigroups $(S_{n}(t))_{t\geq 0}$, i.e., Assumption 3.6(ii). For $s=1$, $(S_{\infty}(t))_{t\geq 0}$ is $L^{\infty}(\mathbb{T})$-contractive since the $L^{1}(\mathbb{T})$-norm of its heat kernel coincides with the $L^{1}(\mathbb{R})$-norm of the Gauss–Weierstrass kernel, which is equal to $1$. For finite $n$, the $L^{\infty}(\mathcal{M}_{n})$-contractivity of $S_{n}(t)=e^{-tA_{n}}$ is equivalent to $A_{n}$ being diagonally dominant with positive diagonal by [64, Lemma 6.1], which holds for Laplacian matrices. Since these assertions can be extended to all $s\in(0,1]$ by subordination, see for instance [50, Theorem 15.2.17], we indeed find that Assumption 3.6(ii) holds. Thus, we can proceed to argue as in b, using Theorem 3.14c and Lemma 3.15b for an arbitrary $\beta\in(\frac{1}{4s},\frac{1}{2})$, to obtain (A1-B)(A4-B) with $\widetilde{M}_{S}=1$ and $\theta+2\beta<1$ for $q\in[2,\infty)$ large enough.
@@ 4. Infinite-dimensional Ornstein–Uhlenbeck process
Here, $A_{n}\colon\mathsf{D}(A_{n})\subseteq E_{n}\to E_{n}$ is a linear operator and $T\in(0,\infty)$ is a time horizon. Moreover, we take $W_{n}\coloneqq\Pi_{n}W_{\infty}$, where $(W_{\infty}(t))_{t\geq 0}$ denotes a cylindrical Wiener process on $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$ taking values in a separable Hilbert space $(H_{\infty},\langle\,\cdot\,,\,\cdot\,\rangle_{H_{\infty}})$ and the operator $\Pi_{n}\in\mathscr{L}(H_{\infty};H_{n})$ is as in assumption (A1) below. Thus, the formal time derivative $\dot{W}_{\infty}$ of $W_{\infty}$ represents space–time Gaussian white noise and $(W_{n})_{t\geq 0}$ is an $H_{n}$-valued $Q_{n}$-cylindrical Wiener process colored in space by the covariance operator $Q_{n}\coloneqq\Pi_{n}\Pi_{n}^{*}\in\mathscr{L}(H_{n})$.
@@
    The operators $(A_{n})_{n\in\overline{\mathbb{N}}}$ on $(E_{n})_{n\in\overline{\mathbb{N}}}$ are uniformly sectorial of angle less than $\frac{1}{2}\pi$ (see Appendix C for the definition of this concept). In particular, their negatives generate bounded analytic $C_{0}$-semigroups $(S_{n}(t))_{t\geq 0}\subseteq\mathscr{L}(E_{n})$ which satisfy Assumption 2.2 with $w=0$. Moreover, there exists a $\beta\in[0,\frac{1}{2})$ such that $R_{n}^{\beta}\coloneqq(\operatorname{Id}_{n}+A_{n})^{-\beta}\in\gamma(H_{n};E_{n})$ for every $n\in\overline{\mathbb{N}}$.
@@
As remarked in Section 2.2, we mainly have applications in mind where the problem corresponding to $n=\infty$ is interpreted as a spatiotemporal stochastic partial differential equation. In the present linear setting, its solution is also known as an infinite-dimensional Ornstein–Uhlenbeck process, and solutions to (4.1) for $n\in\mathbb{N}$ are spatially discretized approximations. Therefore, it is natural to ask whether we can identify the right mode of convergence of the operators $(A_{n})_{n\in\mathbb{N}}$ to $A_{\infty}$ as $n\to\infty$ to ensure the convergence of the processes $(W_{A_{n}})_{n\in\mathbb{N}}$ to $W_{A_{\infty}}$.
@@
Before stating any discrete-to-continuum results, let us introduce some convenient notation for this goal. Using the operators $(\Lambda_{n})_{n\in\mathbb{N}}$ and $(\Pi_{n})_{n\in\mathbb{N}}$, we can take a mapping which has $E_{n}$ as its domain or state space, and turn it into an analogous mapping from or to $E_{\infty}$. For instance, we define the $E_{\infty}$-valued processes $\widetilde{W}_{A_{n}}\coloneqq\Lambda_{n}W_{A_{n}}$, as well as the operators $\widetilde{R}_{n}^{\alpha}\coloneqq\Lambda_{n}R_{n}^{\alpha}\Pi_{n}$ and $\widetilde{S}_{n}(t)\coloneqq\Lambda_{n}S_{n}(t)\Pi_{n}$ in $\mathscr{L}(E_{\infty})$ for $\alpha,t\in[0,\infty)$. Now we can formulate our notion of the convergence $A_{n}\to A_{\infty}$ as $n\to\infty$ as follows:
@@
    For every $x\in E_{\infty}$, we have $\widetilde{R}_{n}^{1}x\to R_{\infty}^{1}x$ in $E_{\infty}$ as $n\to\infty$. Moreover, given $\beta\in[0,\frac{1}{2})$ as in (A2), we have $\widetilde{R}_{n}^{\beta}\to R_{\infty}^{\beta}$ in $\gamma(H_{\infty};E_{\infty})$.
@@
Arguing as in the proof of Proposition B.3 and using (ii), it follows that we only need to prove the claim for all $R$ belonging to some dense subset $D$ of $L^{2}(a,b;\gamma(H;E))$. Note that every $R\in L^{2}(a,b;\gamma(H;E))$ can be approximated by a step function $\sum_{j=1}^{N}\mathbf{1}_{(a_{j},b_{j}^{\prime})}\otimes T_{j}$ with $a<a_{j}^{\prime}<b_{j}^{\prime}<b$ and $T_{j}\in\gamma(H;E)$, and by definition of $\gamma(H;E)$ the latter can be approximated by finite-rank operators. By linearity, it thus suffices to prove the statement for $R$ of the form
@@
Substituting this representation, using (2.5) and (i), we find as $n\to\infty$:
@@
condition (ii) is verified by applying (C.3) to $\|M_{n}(t)\|_{\mathscr{L}(E_{\infty})}$ combined with Assumptions (A1) and (A2) as in (4.16), and hypothesis (i) holds by Lemma C.1. ∎
@@
The proof is based on the Da Prato–Kwapień–Zabczyk factorization method, first formulated in [19] for Hilbert spaces (see also [20, Section 5.3]), and later extended to UMD-type-2 Banach spaces in [9, Theorem 3.2]. The general idea is to express the process $W_{A_{n}}$ as the ‘product’ $\mathfrak{I}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}W^{\frac{1}{2}+\beta^{\prime}}_{A_{n}}$ of a fractional parabolic integral operator $\mathfrak{I}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}$ as in Appendix B and auxiliary process $W^{\frac{1}{2}+\beta^{\prime}}_{A_{n}}$ as in (4.8), and using the smoothing properties of the former.
@@ 5. Approximation of semilinear stochastic evolution equations with additive cylindrical Wiener noise
In Section 5.1 we will assume in particular that the drifts $(F_{n})_{n\in\overline{\mathbb{N}}}$ are uniformly globally Lipschitz and of linear growth to obtain the existence of unique global solutions $(X_{n}(t))_{t\in[0,T]}$, whose lifted counterparts $\widetilde{X}_{n}$ converge to $X_{\infty}$ in $L^{p}(\Omega;C([0,T];E_{\infty}))$ as $n\to\infty$, where $p\in[1,\infty)$ is the stochastic integrability of the initial data. These assumptions are relaxed in Section 5.2, where we suppose that the drifts are uniformly locally Lipschitz and uniformly bounded near zero. In general, this comes at the cost of obtaining merely local solutions, converging in a weaker norm. However, if one can show independently that the solutions are global and the $L^{p}(\Omega;C([0,T];E_{n}))$-norms of $X_{n}$ are uniformly bounded in $n\in\mathbb{N}$, then we recover the stronger sense of convergence.
@@ 5.1. Globally Lipschitz drifts of linear growth
$\widetilde{\Phi}_{n,T}(u)=\widetilde{S}_{n}\otimes\widetilde{\xi}_{n}+\widetilde{S}_{n}*\widetilde{F}_{n}(\,\cdot\,,u)+\widetilde{W}_{A_{n}}$ (5.15)
@@
$\|X_{\infty}-\widetilde{X}_{n}\|_{L^{p}(\Omega;C([0,\frac{3}{2}T_{0}];E_{\infty}))}$ $\leq\|X_{\infty}-\widetilde{X}_{n}\|_{L^{p}(\Omega;C([0,\frac{1}{2}T_{0}];E_{\infty}))}$ (5.34)
@@ 6.1. Setting and convergence for globally Lipschitz drifts
Fix a $\beta^{\prime}\in(\beta,\frac{1}{2})$ such that $\theta+2\beta^{\prime}<1$, where $\beta\in[0,\frac{1}{2})$ is as in (A2). Without loss of generality, we may assume that $p\in(2,\infty)$ is so large that, in fact, $\theta+2\beta^{\prime}<1-\frac{2}{p}$. As in the proof of Proposition 4.5, we see that $W_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}(\omega,\,\cdot\,)$, where $W_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}$ is the auxiliary process defined by (4.8), belongs to $L^{p}(0,T;E_{n})$ for $\mathbb{P}$-a.e. $\omega\in\Omega$, hence $\mathfrak{I}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}W_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}(\omega,\,\cdot\,)\in C([0,T];B_{n})$ by applying Proposition B.1b with $E\coloneqq E_{n}$, $F\coloneqq B_{n}$ and $\alpha\coloneqq\theta$ from assumption (A4-B). Moreover, one finds that the process $\mathfrak{I}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}W_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}$ is a continuous modification of $W_{A_{n}}$, belonging to $L^{p}(\Omega;C([0,T];B_{n}))$. It now suffices to establish the following, as $n\to\infty$: